# HyGen: Efficient LLM Serving via Elastic Online-Offline Request Co-location

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have facilitated a wide range of applications with distinct quality-of-experience requirements, from latency-sensitive online tasks, such as interactive chatbots, to throughput-focused offline tasks like document summarization. While deploying dedicated machines for these services ensures high-quality performance, it often results in resource underutilization.
This paper introduces HyGen, an interference-aware LLM serving system that enables efficient co-location of online and offline workloads while preserving latency requirements. HyGen incorporates two key innovations: (1) performance control mechanisms, including a latency predictor for batch execution time estimation and an SLO-aware profiler to quantify interference, and (2) SLO-aware offline scheduling policies that maximize throughput and prevent starvation, without compromising online serving latency. Our evaluation on production workloads shows that HyGen achieves up to 5.84$\times$ higher throughput compared to existing advances while maintaining comparable latency. The code will be released soon.
@@ 1 Introduction
Large language models (LLMs) have emerged as transformative tools across diverse domains, powering applications from interactive chatbots to document analysis systems Grand View Research (2023). As LLM services continue to spread, their serving deployment must handle both latency-critical online requests (e.g., chatbot interactions OpenAI (2022); Liu et al. (2024a)) and throughput-oriented offline tasks (e.g., document summarization Grand View Research (2023)). Online serving demands low and stable response times, measured by Time to First Token (TTFT) and Time Between Tokens (TBT), to ensure smooth user interactions. In contrast, offline tasks prioritize high throughput and resource utilization, often processing large batches of requests with relaxed latency constraints. The disparity in these requirements has led most production deployments to segregate online and offline serving onto separate GPU clusters Sheng et al. (2023); Zhao et al. (2024); Patel et al. (2024); Stojkovic et al. (2024).
However, real-world LLM workloads demonstrate significant temporal variations in request loads. Our analysis of production traces shows that in addition to the popular diurnal request arrival patterns, online request rates can vary by up to 3$\times$ within minutes (Section 3). To maintain acceptable latency under such bursty loads, service providers must provision GPU resources for peak demand, leading to substantial underutilization during off-peak hours.
This resource overprovisioning suggests an opportunity to improve resource efficiency via hybrid serving—co-locating online and offline workloads on the same GPU cluster. By opportunistically co-locating offline tasks with online tasks during periods of low online load, the system could maintain high GPU utilization while preserving latency guarantees for online requests. However, realizing this opportunity requires addressing several fundamental challenges.
@@
Second, LLM workloads are inherently unpredictable in both their arrival patterns and resource demands. Online requests arrive in unexpected bursts, while offline tasks appear in irregular batch intervals. This variability is further complicated by uncertainty in resource demands—input sequences vary widely in length, and the number of output tokens can hardly be predicted until generation completes.
Third, co-locating online and offline workloads introduces interference, which degrades the performance of both. For instance, offline requests can delay time-sensitive online requests, while using large batch sizes to boost offline request throughput can negatively impact the latency of online requests. Effectively managing this interference while ensuring performance guarantees demands meticulous orchestration of resource sharing.
This paper presents HyGen, an interference-aware LLM serving system that elastically co-locates online and offline workloads. HyGen introduces several key techniques: (1) a latency predictor that accurately estimates the execution time of heterogeneous request batches, (2) an interference-aware profiler that quantifies the performance impact of co-location, and (3) an adaptive scheduler that maximizes offline throughput while maintaining strict latency guarantees for online serving.
Our evaluation on production workloads shows that HyGen guarantees latency requirements in co-locating requests and achieves up to 5.84$\times$ higher throughput compared to existing advances while maintaining comparable latency.
To summarize, this paper makes the following contributions:
@@
  • We propose HyGen, a new serving system that enables efficient resource sharing, improving throughput with guaranteed performance isolation.
  • To improve the performance isolation and throughput of HyGen, we design performance control mechanisms and SLO-aware offline scheduling policies.
@@ 2 Related Work
Recent works have improved GPU computation through optimized kernels Dao et al. (2022); Ye et al. (2025), memory efficiency through efficient KV-cache management techniques Lin et al. (2024a); Liu et al. (2024b); Qin et al. (2024), distributed inference through various model parallelism strategies Shoeybi et al. (2019); Pope et al. (2023); Liu et al. (2023), and model compression through quantization approaches Xiao et al. (2023); Lin et al. (2024b); Dettmers et al. (2022). Orca Yu et al. (2022) introduces iteration-level scheduling to eliminate padding overhead in variable-length sequence batching, and enables early exits for completed requests, thereby reducing batching latency. vLLM Kwon et al. (2023) develops PagedAttention for efficient memory management, while Sarathi-Serve proposes chunked prefill to minimize inter-request interference within batches. However, few existing systems explicitly address the challenge of co-locating latency-critical online requests with offline workloads to maximize resource utilization—a key focus of our work.
@@
As offline inference gains increasing traction, Huggingface Accelerate Hugging Face (2025), DeepSpeed ZeRO-Inference Aminabadi et al. (2022) and FlexGen Sheng et al. (2023) enable efficient LLM serving on commodity GPUs by offloading strategies. FlexGen aggregates memory from GPU, CPU and disk to maximize serving throughput. For KV cache optimization, InfiniGen Lee et al. (2024) proposes dynamic cache management through selective prefetching based on token importance prediction. S3 Jin et al. (2023) tackles the KV cache problem through output sequence length prediction, allowing more efficient memory allocation by avoiding over-allocation of the maximum sequence length. NanoFlow Zhu et al. (2024) focuses on maximizing hardware utilization through intra-device parallelism, introducing nano-batching and execution unit scheduling to overlap compute-, memory-, and network-intensive operations.
@@
Co-locating latency-critical applications with batch applications is a widely adopted approach to improve resource utilization in datacenters, with the popular resource partitioning techniques Chen et al. (2019b); Zhu & Erez (2016); Lo et al. (2015); Wu & Martonosi (2008); Chen et al. (2019a); Kasture et al. (2015); Sanchez & Kozyrakis (2011); Kasture & Sanchez (2014); Zhang et al. (2019); Iorgulescu et al. (2018); Patel & Tiwari (2020); Chen et al. (2023). Prior work has also explored multiplexing to serve multiple DNNs Dhakal et al. (2020); Tan et al. (2021); Zhang et al. (2023b), LLMs Duan et al. (2024), or LoRAs Chen et al. (2024); Sheng et al. (2024a); Wu et al. (2024). However, prior efforts have largely overlooked the opportunity to co-locate online and offline inference requests within the same LLM.
@@ 3.1 LLM Serving
LLM generation consists of two sequential stages: (i) Prefilling Stage: It processes the entire input prompt in parallel to generate the first output token, computing and caching intermediate model representations (i.e., the KV cache) of the input sequence to avoid regeneration in subsequent tokens, and then the (ii) Decoding Stage: the model generates tokens one at a time in an auto-regressive manner. Each iteration processes only the most recent token while utilizing the KV cache of previous tokens, reducing computational overhead. This stage iterates until meeting a termination condition, such as generating an end-of-sequence token or reaching the maximum length.
Model executions in the prefilling stage are compute-intensive due to the high degree of parallelism in processing the entire input sequence simultaneously. Yet in the decoding phase, the same operations become memory-intensive due to the reliance on the KV cache Agrawal et al. (2024), and the total memory access cost grows linearly with the sequence length as each new token’s KV cache is appended.
@@
The second counterpart is offline serving, which prioritizes throughput over latency, measured in requests per second (RPS) or tokens per second (TPS) for prefill/decode operations. Key applications include model benchmarking Hendrycks et al. (2021b); Meta-Team (2024), document processing Chen et al. (2021); Zhang et al. (2023a), data cleaning Narayan et al. (2022); Zheng et al. (2023), and synthetic data generation Allal et al. (2024).
@@ 3.2 Motivation and Challenges
(a) 1 hour
(b) 2 minutes
Figure 1: Request rate variation in Microsoft Azure’s LLM service over one-hour and two-minute periods, showing significant temporal fluctuations between peak and off-peak periods.
@@
This load variability suggests an opportunity to improve GPU utilization by co-locating online inference requests with offline batch workloads. During periods of low online traffic, the system could opportunistically schedule offline tasks to harvest idle GPU resources. However, co-locating these workloads introduces several fundamental challenges:
@@
    Diverse Latency Requirements: Applications and even requests of an application have distinct latency requirements. For example, users of different ages have different reading speeds, thus TBT requirements Liu et al. (2024a). These requirements are often expressed through multiple statistical metrics - some services prioritize tail latency guarantees (e.g., P99 TTFT $\leq$ 1000ms), others focus on average performance (e.g., mean TBT $\leq$ 50ms), and many need to satisfy both simultaneously. How to provide latency guarantees for diverse requests while sharing resources?
@@
    Massive Uncertainties: LLM serving faces temporal uncertainty—online requests arrive in unpredictable bursts with varying urgency levels—as well as resource demand uncertainty due to unpredictable output lengths. A seemingly simple query might trigger an elaborate response requiring thousands of tokens. How to perform efficient scheduling in the wild?
@@
    Request Interference: Co-locating online and offline workloads introduces interference that degrades system performance. Large batches of offline requests arriving simultaneously can cause severe head-of-line blocking, delaying the processing of time-sensitive online requests. Worse, batching requests of long inputs (e.g., a long document summarization request with 2,048 tokens) with short interactive queries (e.g., 128 tokens) can result in a dramatic increase in attention computation overhead, elongating the latency for all requests in the batch by an order of magnitude Agrawal et al. (2024). How to account for interference in co-locating requests?
@@ 4.1 Overview
Figure 2: HyGen overview: Online requests and offline requests sent by users are asynchronously selected and processed.
@@
3:      latency budget $t$, remaining chunk size $c$,
4:      memory budget $m$
5:     Output: batched requests $B$
6:     $B\leftarrow\{\}$
7:     for $r\in R$.decoding do
8:        $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
9:        if $t_{req}\leq t$ then
10:           $t\leftarrow t-t_{req}$
11:           $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
12:        end if
13:     end for
14:     for $r\in R$.prefilling $\cup Q$ do
15:        TRY_SCHEDULE:
16:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
17:        if $l>0$ then
18:           $t\leftarrow t-t_{req}$
19:           $c\leftarrow c-l$
20:           $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
21:           $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
22:        else
23:           if PHASE ==absent$=$ = ONLINE and $R\neq B$ then
24:              PERFORM_PREEMPTION($R$, $m$)
25:              goto TRY_SCHEDULE
26:           else
27:              break
@@
29:        end if
30:     end for
31:     return $B$
32:  end function
Figure 2 illustrates the system architecture of HyGen. To handle diverse latency requirements and unpredictable workloads, HyGen maintains separate queues for online and offline requests, each managed by a dedicated request selector that tracks active requests and their states. This separation allows HyGen to support scheduling policies for different request types without the complexity of managing mixed priorities in a single queue.
The scheduling process in HyGen operates in two consecutive phases to carefully manage interference between online and offline requests. In the online phase, the scheduler processes the online request queue to construct an initial batch based on latency SLOs. This phase follows traditional continuous batching principles and supports various scheduling policies such as First-Come-First-Serve (FCFS) Kwon et al. (2023) and Virtual Token Counter (VTC) Sheng et al. (2024b). The scheduler prioritizes running requests before processing the waiting queue, continuing until reaching memory, batch size, or chunk size limits. A detailed workflow of this process is displayed in Algorithm 1. During a default two-phase scheduling, this function is called two times separately for online and offline requests.
During online batch establishment, when the current online batch reaches the memory limit and more online requests remain to be scheduled, HyGen performs a priority-based preemption to maximize SLOs. To handle dynamic request arrivals while maintaining service guarantees, we implement an online-first preemption mechanism, where online requests are prioritized over all offline requests to minimize interference. Offline requests are prioritized based on their original priority in the waiting queue and their memory consumption. Preempted offline requests are removed from the current running list and reinserted into the offline queue, where they maintain their execution state to enable efficient rescheduling in subsequent batches.
After establishing the online batch, the scheduler enters the offline phase where it opportunistically fills the remaining capacity with offline requests. By utilizing a latency predictor to estimate batch execution time and determine remaining latency budget, instead of relying on static chunk size or batch size limits, we achieve both higher accuracy in interference prediction and strong transferability across different hardware configurations and workload patterns. This enables higher throughput while maintaining SLO guarantees by eliminating the need for conservative scheduling. The selection of offline requests then considers this available batch execution latency budget derived from user-specified SLOs, opportunities for efficient execution through prefix sharing, and current batch composition and resource constraints. Within this phase, running requests are prioritized over waiting ones to maintain execution efficiency.
HyGen addresses these complex scheduling decisions through two key components detailed in the following sections. Section 4.2 describes our performance control mechanisms, including the latency predictor and SLO-aware profiler that enable accurate resource allocation decisions. Section 4.3 presents our SLO-aware offline scheduling policies that maximize throughput while maintaining performance isolation through careful workload co-location. A detailed algorithm for the entire HyGen scheduling workflow can be found in Appendix A.1.
@@ 4.2 Performance Control Mechanisms
Efficient co-location of online and offline requests requires precise control over execution timing and resource allocation. The key challenge lies in accurately estimating the performance impact of scheduling decisions while ensuring end-to-end SLO compliance. This section presents two core mechanisms in HyGen that address this challenge: a latency predictor for estimating batch execution times and an SLO-aware profiler for translating these estimates into scheduling decisions.
@@
The design of our latency predictor is guided by three key requirements. First, it must provide fast inference to support real-time scheduling decisions when evaluating different batch compositions. Second, it needs to be robust across varying workload patterns conditions to maintain reliable performance. Third, it should be adaptable to different hardware configurations to capture their unique performance characteristics.
The execution time of an LLM serving batch is primarily influenced by two distinct processing stages with different computational patterns. The prefill stage exhibits quadratic complexity due to attention computations, with latency growing quadratically with input sequence length. This stage’s total load depends on both the number of requests and their individual sequence lengths. In contrast, the decode stage shows linear scaling, with computational requirements growing proportionally with the number of tokens.
To capture these characteristics, we model batch execution time as:
@@
We implement this model using linear regression, chosen for its combination of efficiency and effectiveness. The linear model enables rapid evaluation of different batch compositions during scheduling, while its simple feature set ensures stable predictions across varying conditions. The training process involves three steps:
  1. 1.
    Training Data Collection: We systematically profile target hardware across diverse batch compositions, varying the number of requests in different phases, sequence length distributions, and total batch sizes.
  2. 2.
    Feature Engineering: For each batch, we compute six key features ($S_{p}$, $S_{d}$, $S_{p}^{2}$, $S_{d}^{2}$, $N_{p}$, $N_{d}$) based on token counts and batch sizes, capturing both linear and quadratic scaling.
  3. 3.
    Model Training: Linear regression coefficients are optimized to minimize prediction error across the training set, providing sufficient accuracy for scheduling decisions while maintaining fast inference times.
@@
While accurate latency prediction is essential, translating these predictions into scheduling decisions requires careful consideration of end-to-end performance requirements. Our SLO-aware profiler addresses this by determining appropriate latency budgets that ensure SLO compliance under various operating conditions.
The profiler first analyzes specific workload-SLO combinations to establish viable latency budget ranges. Then, given that SLO has a positive correlation with batch latency, we perform the binary search on the batch latency budget to get an upper limit that meets the SLO. This profiling enables three key capabilities. First, it determines appropriate latency thresholds that maintain SLO compliance for different workload types. Second, it provides flexible adaptation by adjusting budgets based on changing workload characteristics and performance requirements. Third, it establishes a robust foundation for hybrid scheduling by accounting for both online and offline workload patterns.
Together, these mechanisms enable HyGen to make precise scheduling decisions that maximize resource utilization while maintaining strict performance guarantees. The latency predictor provides the quantitative basis for evaluating scheduling options, while the SLO-aware profiler ensures these decisions translate into consistent end-to-end performance.

### v2

@@ Abstract
Large language models (LLMs) have facilitated a wide range of applications with distinct service-level objectives (SLOs), from latency-sensitive online tasks like interactive chatbots to throughput-oriented offline workloads like document summarization. The existing deployment model, which dedicates machines to each workload, simplifies SLO management but often leads to poor resource utilization.
This paper introduces HyGen, an interference-aware LLM serving system that enables efficient co-location of online and offline workloads while preserving latency requirements. HyGen incorporates two key innovations: (1) performance control mechanisms, including a latency predictor to estimate batch execution time and an SLO-aware profiler to quantify latency interference, and (2) SLO-aware offline scheduling policies that maximize serving throughput and prevent starvation, without compromising online serving latency. Our evaluation on production workloads shows that HyGen achieves up to 3.87$\times$ overall throughput and 5.84$\times$ offline throughput gains over online and hybrid serving baselines, respectively, while strictly satisfying latency SLOs.
@@ 1 Introduction
Large language models (LLMs) have emerged as transformative tools across diverse domains, handling both latency-critical online requests (e.g., chatbot interactions OpenAI (2022); Liu et al. (2024a)) and throughput-oriented offline tasks (e.g., document summarization Grand View Research (2023)). Online serving demands low and stable response times, measured by Time to First Token (TTFT) and Time Between Tokens (TBT), while offline tasks prioritize high throughput and resource utilization, often processing large batches with relaxed latency constraints. The disparity in these requirements has led most production deployments to segregate online and offline serving onto separate GPU clusters Sheng et al. (2023); Zhao et al. (2024); Patel et al. (2024); Stojkovic et al. (2024).
However, real-world LLM workloads demonstrate significant temporal variations in request load. Our analysis of production traces shows that in addition to the popular diurnal request arrival patterns, online request rates can vary by up to 3$\times$ within minutes (Section 3). To meet latency requirements under such bursty loads, service providers must provision GPU resources for peak demand, leading to substantial underutilization during off-peak hours.
This resource overprovisioning suggests an opportunity to improve resource efficiency via hybrid serving—co-locating online and offline workloads on the same GPU cluster. By opportunistically padding online requests with offline requests during periods of low online load, the system could maintain high GPU utilization while preserving latency guarantees for online requests. However, realizing this opportunity requires addressing several fundamental challenges.
@@
Second, LLM workloads are inherently unpredictable in both their arrival patterns and resource demands. Request arrival rates exhibit both diurnal patterns and unpredictable short-term fluctuations. This variability is further complicated by uncertainty in resource demands—input sequences vary widely in length, and the number of output tokens can hardly be predicted until generation completes.
Third, co-locating online and offline workloads introduces interference. For instance, offline requests may delay time-sensitive online requests; using large batch sizes to improve offline request throughput can increase the latency of online requests. Effectively managing this interference while ensuring performance guarantees demands meticulous orchestration of resource sharing.
This paper presents HyGen, an interference-aware LLM serving system that elastically co-locates online and offline workloads. HyGen introduces several key techniques: (1) a latency predictor that accurately estimates the execution time of different request batches, (2) an interference-aware profiler that quantifies the performance interference of co-location, and (3) an adaptive scheduler that maximizes offline throughput while maintaining strict latency guarantees for online requests.
Our evaluation on production workloads shows that HyGen improves serving throughput by 3.87-5.84$\times$ over existing advances Agrawal et al. (2024), while guaranteeing strict SLO compliance. To summarize, this paper makes the following contributions:
@@
  • We propose HyGen, a new serving system that enables efficient resource sharing, improving throughput with performance isolation.
@@ 2 Related Work
Recent works have improved GPU computation through optimized kernels Dao et al. (2022); Ye et al. (2025), memory efficiency through efficient KV-cache management techniques Lin et al. (2024a); Liu et al. (2024b); Qin et al. (2024), distributed inference through various model parallelism strategies Shoeybi et al. (2019); Pope et al. (2023); Liu et al. (2023), and model compression approaches Xiao et al. (2023); Lin et al. (2024b); Dettmers et al. (2022). Orca Yu et al. (2022) introduces iteration-level scheduling to eliminate padding overhead in variable-length sequence batching and enables early exits for completed requests. vLLM Kwon et al. (2023) develops PagedAttention for efficient memory management, while Sarathi-Serve Agrawal et al. (2024) proposes chunked prefill to minimize inter-request interference within batches. However, few existing systems address the challenge of co-locating latency-critical online requests with offline workloads.
@@
As offline inference gains increasing traction, Huggingface Accelerate Hugging Face (2025), DeepSpeed ZeRO-Inference Aminabadi et al. (2022) and FlexGen Sheng et al. (2023) enable efficient LLM serving on commodity GPUs by offloading strategies. FlexGen aggregates memory from GPU, CPU and disk to maximize serving throughput. S3 Jin et al. (2023) utilizes output sequence length prediction technique, allowing more efficient memory allocation by avoiding over-allocation of the maximum sequence length.
@@
Co-locating latency-sensitive applications with batch applications has been widely adopted to improve resource utilization in datacenters Chen et al. (2019); Zhu & Erez (2016); Zhang et al. (2019); Iorgulescu et al. (2018); Patel & Tiwari (2020); Chen et al. (2023). Prior work has also explored resource multiplexing to serve multiple DNNs Dhakal et al. (2020); Tan et al. (2021); Zhang et al. (2023b), LLMs Duan et al. (2024), or LoRAs Chen et al. (2024); Sheng et al. (2024a); Wu et al. (2024). However, prior efforts have largely overlooked the opportunity to co-locate online and offline inference requests within the same LLM.
@@ 3.1 LLM Serving
LLM generation consists of two sequential stages: (1) Prefill Stage: It processes the input prompt in parallel to generate the first output token, computing and caching intermediate model representations (i.e., the KV cache) to avoid regeneration in subsequent tokens, and the (2) Decode Stage: the model generates tokens one at a time in an auto-regressive manner. Each iteration processes only the most recent token while utilizing the KV cache of previous tokens, reducing computational overhead.
Model executions in the prefill stage are compute-intensive due to the high degree of parallelism in processing the entire input sequence simultaneously. Yet in the decode phase, the same operations become memory-intensive due to the reliance on the KV cache Agrawal et al. (2024), and the total memory access cost grows linearly with the sequence length as each new token’s KV cache is appended.
@@
The second scenario is offline serving, which prioritizes throughput over latency, measured in queries per second (QPS) or tokens per second (TPS). For example, OpenAI’s Batch API processes requests with relaxed latency requirements (up to 24 hours) at significantly lower costs compared to standard APIs OpenAI (2024). Applications include model benchmarking Hendrycks et al. (2021b); Meta-Team (2024), document processing Chen et al. (2021); Zhang et al. (2023a), data cleaning Narayan et al. (2022); Zheng et al. (2023), and synthetic data generation Allal et al. (2024).
@@ 3.2 Motivation and Challenges
(a) One-hour Request Rates
(b) Two-minute Request Rates
Figure 1: Request rate varies significantly in Microsoft Azure’s LLM service over one-hour and two-minute periods.
@@
This load variability suggests an opportunity to improve GPU utilization by co-locating online requests with offline requests. During periods of low online traffic, the system could opportunistically schedule offline tasks to harvest idle GPU resources. While recent methodologies in simultaneous batching of prefill and decode requests set the premise for dynamic request co-location Agrawal et al. (2024), doing so at scale introduces several fundamental challenges:
@@
    Diverse Latency Requirements: Applications and even requests of an application have distinct latency requirements. For example, paid users require strict latency SLOs while free users accept more relaxed guarantees. How to flexiblely specify latency guarantees for online requests when sharing resources?
@@
    Massive Uncertainties: LLM serving faces temporal uncertainty—online requests arrive in unpredictable bursts with varying urgency levels—as well as resource demand uncertainty due to unpredictable output lengths. How to perform efficient scheduling in the wild?
@@
    Request Interference: Co-locating online and offline workloads introduces performance interference. Large batches of offline request can cause severe head-of-line blocking, delaying the processing of time-sensitive online requests. Worse, batching requests of long inputs with short interactive queries can elongate the latency of all requests in the batch by an order of magnitude Agrawal et al. (2024). How to account for interference in co-locating requests?
@@ 4.1 Overview
Figure 2: HyGen overview. Online requests and offline requests are asynchronously processed. Offline requests are opportunistically scheduled while retaining the batch latency budget.
@@
3:      latency budget $t$, chunk size $c$, memory budget $m$
4:     Output: batched requests $B$
5:     $B\leftarrow\{\}$
6:     for $r\in R$.decode do
7:        $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
8:        if $t_{req}\leq t$ or PHASE ==absent$=$ = ONLINE then
9:           $t\leftarrow t-t_{req}$
10:           $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
11:        end if
12:     end for
13:     for $r\in R$.prefill $\cup Q$ do
14:        TRY_SCHEDULE:
15:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
16:        if $l>0$ then
17:           $t\leftarrow t-t_{req}$
18:           $c\leftarrow c-l$
19:           $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
20:           $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
21:        else
22:           if PHASE ==absent$=$ = ONLINE and $R\neq B$ then
23:              PERFORM_PREEMPTION($R$, $m$)
24:              goto TRY_SCHEDULE
25:           else
26:              break
27:           end if
@@
29:     end for
30:     return $B$
31:  end function
Figure 2 illustrates the overall design of HyGen. To handle diverse SLOs and unpredictable workloads, HyGen maintains separate queues for online and offline requests, each managed by a dedicated request selector. This design makes HyGen compatible with various existing scheduling policies for different request types. Following Sarathi-Serve Agrawal et al. (2024), we employ stall-free batching and chunked-prefills, batching prefill and decode requests together for efficient utilization of compute resources and dynamic co-location of requests.
HyGen’s scheduling process occurs in two phases. The first (online) phase handles the online request queue to construct an initial batch, supporting various scheduling policies such as First-Come-First-Serve (FCFS) Kwon et al. (2023) and Virtual Token Counter (VTC) Sheng et al. (2024b). If the memory limit is reached and more online requests remain, a priority-based preemption mechanism is employed by prioritizing online requests over offline requests to minimize interference. Preempted offline requests are reclaimed into the offline queue, where they maintain their execution state to enable efficient rescheduling in subsequent batches.
In the second (offline) phase, the remaining capacity is filled with offline requests using a latency predictor to capture the latency impact of request filling. The predictor calculates the latency of the online batch and selects offline requests based on the available batch execution latency budget, i.e., the latency slack without violating the latency requirement of online requests. It predicts the latency cost for decode requests, and calculates a maximum chunked prefill length for prefill requests for the given latency budget. Offline requests are scheduled in order until reaching the resource capacity, based on our SLO-aware scheduling policy. A formalized algorithm of the SLO-aware scheduling process is displayed in Algorithm 1. Appendix A.1 further presents a holistic workflow of two-phase scheduling, including the management of online and offline request queues and implementation of our SLO-aware scheduling algorithm.
Section 4.2 describes our performance control mechanisms, including the latency predictor and SLO-aware profiler that enable accurate resource allocation. Section 4.3 presents our SLO-aware offline scheduling policies that maximize throughput while maintaining SLOs.
@@ 4.2 Performance Control Mechanisms
Co-locating online and offline requests requires precise control over execution timing and resource allocation. The key challenge lies in accurately estimating the latency impact of scheduling decisions to ensure end-to-end SLO compliance. This section introduces a latency predictor for estimating batch execution latency and an SLO-aware profiler for translating the estimates into scheduling decisions.
@@
The design of our latency predictor is guided by three key requirements. First, it must provide fast inference to support real-time scheduling decisions. Second, it needs to be robust across varying workload patterns to maintain reliable performance. Third, it should be adaptable to different hardware configurations to accurately capture their unique performance characteristics.
The execution time of an LLM serving batch is primarily determined by two distinct processing stages with different computational patterns. The prefill stage exhibits quadratic complexity due to attention computations, with latency growing quadratically with input sequence length. The total load of this stage depends on both the number of requests and their individual sequence lengths. In contrast, the decode stage shows linear scaling, with computational requirements growing proportionally with the number of tokens. We can model the batch execution time below to capture these characteristics:
@@
We employ linear regression as the prediction model because of its efficiency and effectiveness. Training data for the model is collected by systematically profiling target hardware across diverse batch compositions, varying in the number of requests in different phases, sequence length distributions, and total batch sizes. The linear model enables rapid evaluation of different batch compositions during scheduling, while its simple feature set ensures stable predictions across varying conditions.
@@
The latency predictor provides accurate latency estimates for filling offline requests. Our SLO-aware profiler leverages a latency budget to ensure SLO compliance in scheduling.
The profiler first analyzes the given combination of workload and SLO to establish viable latency budget ranges. Given that a larger batch size and longer inputs will increase batch latency, the profiler test-runs latency budgets within the range to check their compliance with the given SLO and employs binary search to decide an upper limit that meets the overall SLO for online requests. During deployment, this latency budget is then used as the batch latency limit in the two-phase SLO-aware scheduling process (Section 4.1) to ensure SLO compliance. This profiling enables three key capabilities. First, it determines appropriate latency thresholds that maintain SLO compliance for different workload types. Second, it provides flexible adaptation by adjusting budgets based on changing workload characteristics and performance requirements. Third, it establishes a robust foundation for hybrid scheduling by accounting for both online and offline workload patterns.